Explain how to calculate the volumes of the following shapes:

• a cuboid

• a sphere

• a cone

• a triangular prism

• a cylinder


1
Expert's answer
2021-05-17T16:08:03-0400

Cuboid whose length = l, width = w and height = h

To find the volume of a cuboid, multiply its length by its width by its height. This can be written as: v o l u m e = l × w × h.


Sphere whose radius = "r"

Volume of sphere ="\dfrac{4}{3}\pi r^3"


Cone having height = "h" , radius = "r" and slant height = "l"

Volume of cone ="\dfrac{1}{3}\pi r^2h"


The general formula for calculating the volume of any prism is:

Volume = (Area of the base) "\times" length

For a triangular prism, the formula would be:

Volume = (Area of the triangular base) "\times" length 


Cylinder whose radius = "r" , height = "h"

Volume of cylinder ="\pi r^2 h"